Bibliography

proceeding

23/03-04 16:20

[acm1991a] ACM. Communications Architectures & Protocols, SIGCOMM’91 Conference Proceedings, volume 21(4) of ACM Computer Communication Review, Zürich, Switzerland, September 1991. ACM Press.

[acm1991b] ACM. The 18th Annual International Symposium on Computer Architecture, Conference Proceedings, volume 19(3) of Computer Architecture News, Toronto, Canada, May 1991. ACM Press.

[acm1991c] ACM. Fourth International Conference on Architectural Support for Programming Languages and Operating Systems, ASPLOS-IV Proceedings, volume 25 of Operating Systems Review, Santa Clara, California, April 1991. Special Issue.

[acm1991d] ACM. Proceedings of the Thirteenth ACM Symposium on Operating Systems Principle, volume 25(5) of Operating Systems Review, Pacific Grove, California, October 1991. ACM Press.

[acm1992a] ACM. Communications Architectures & Protocols, SIGCOMM’92 Conference Proceedings, volume 22(4) of ACM Computer Communication Review, Baltimore, Maryland, USA, August 1992. ACM Press.

[acm1992b] ACM. The 19th Annual International Symposium on Computer Architecture, Conference Proceedings, volume 20(2) of Computer Architecture News, Gold Coast, Australia, May 1992. ACM Press.

[acm1993a] ACM. Proceedings of the Fourteenth ACM Symposium on Operating Systems Principle, volume 27(5) of Operating Systems Review, Asheville, NC, December 1993. ACM Press.

[acm1994a] ACM. Sixth International Conference on Architectural Support for Programming Languages and Operating Systems, ASPLOS-VI Proceedings, volume 28(5) of Operating Systems Review, San Jose, California, October 1994.

[acm1994b] ACM. Proceedings of the First USENIX Symposium on Operating Systems Design and Implementation (OSDI), volume 28 of Operating Systems Review, Monterey, California, November 1994. ACM Press. Special Issue.

[acm1995a] ACM. Proceedings Multimedia’95, San Francisco, November 1995. ACM Press.

[acm1995b] ACM. Proceedings of the Fifteenth ACM Symposium on Operating Systems Principle, volume 29(5) of Operating Systems Review, Copper Mountain Resort, Colorado, December 1995. ACM Press.

[acm1997a] ACM. Proceedings Multimedia’97, Seattle, Washington, USA, November 1997. Addison-Wesley/ACM Press.

[acm1998a] ACM. Proceedings Multimedia’98, Bristol, England, September 1998. Addison-Wesley/ACM Press.

[davies1998a] Nigel Davies, Kerry Raymond, and Jochen Seitz, editors. Middleware’98, IFIP International Conference on Distributed Systems Platforms and Open Distributed Processing. IFIP, Springer-Verlag, 1998.

[ecoop-aop1997a] Cristina Lopes, Kim Mens, Bedir Tekinerdogan, and Gregor Kiczales, editors. Aspect-Oriented Programming Workshop at ECOOP’97, June 1997.

[ecoop-wcop1997a] Wolfgang Weck, Jan Bosch, and Clemens Szyperski, editors. Proceedings of the Second International Workshop on Component-Oriented Programming (WCOP’97) at ECOOP’97, number 5 in TUCS General Publication. Turku Centre for Computing Science, September 1997.

[europen1991a] EurOpen. Distributed Open Systems in Persective, EurOpen Spring 1991 Conference Proceeding, Tromsø, Norway, May 1991.

[europen1992a] EurOpen and UniForum. Distributed Computing — Practice and Experience, Proceedings of the Autumn 1992 OpenForum Technical Conference, Utrecht, The Netherlands, November 1992.

[ieee-cs1995a] IEEE Computer Society. Proceedings of the International Conference on Multimedia and Networking, Aizu, Japan, September 1995. IEEE Computer Society Press.

[ifip1997a] IFIP. IWQOS’97 Proceedings, Fifth International Workshop on Quality of Service, New York, USA, May 1997.

[lampson1981a] B. W. Lampson, M. Paul, and H. J. Siegert, editors. Distributed Systems — Architecture and implementation, An Advanced Course. Springer-Verlag, 1981.

[nik1992a] Norsk Informatikk Konferanse, Tromsø, November 1992. Tapir.

[nik1993a] Norsk Informatikk Konferanse, Halden, November 1993. Tapir.

[nik1994a] Norsk Informatikk Konferanse, Molde, November 1994. Tapir.

[plagemann1998a] Thomas Plagemann and Vera Goebel, editors. IDMS’98, Proceedings of the 5th International Workshop on Interactive Distributed Multimedia Systems and Telecommunication Services, volume 1483 of Lecture Notes in Computer Science, Oslo, Norway, September 1998. Springer-Verlag.

[python1997a] The Sixth International Python Conference, San Jose, California, October 1997.

[python1998a] The Seventh International Python Conference, Houston, Texas, November 1998.

[python2000a] The Eighth International Python Conference, Arlington, Virginia, January 2000.

[rolia1997a] Jerome Rolia, Jacob Slonim, and John Botsford, editors. Open Distributed Processing and Distributed Platforms, Toronto, Canada, May 1997. IFIP, Chapman & Hall.

[sventek2000a] Joseph Sventek and Geoffrey Coulson, editors. Middleware 2000, IFIP/ACM International Conference on Distributed Systems Platforms, volume 1795 of Lecture Notes in Computer Science, New York, NY, USA, April 2000. IFIP/ACM, Springer-Verlag.

[treleaven1986a] P. Treleaven and M. Vanneschi, editors. Future Parallel Computers, An Advanced Course, volume 272 of Lecture Notes in Computer Science, Pisa, Italy, June 1986. Springer-Verlag.

[troya1997b] José M. Troya and Antonio Vallecillo. On the addition of properties to components. In Wolfgang Weck, Jan Bosch, and Clemens Szyperski, editors, Proceedings of the Second International Workshop on Component-Oriented Programming (WCOP’97), number 5 in TUCS General Publication, pages 95–103. Turku Centre for Computing Science, September 1997.

[ventre1996a] G. Ventre, J. Domingo-Pascual, and A. Danthine, editors. Multimedia Telecommunications and Applications, Third International COST 237 Workshop, volume 1185 of Lecture Notes in Computer Science, Barcelona, Spain, November 1996. Springer-Verlag.